OpenStudy (anonymous):

Find the x-intercept for the linear equation: 4x – 3y = 12 A) (0, -4) B) (0, 3) C) (-4, 0) D) (3, 0) i must be doing something wrong because i keep getting D as my answer and its A according to the answer key . can someone help me step by step?

HanAkoSolo (jamierox4ev3r):

4x-3y=12 -4x -4x -3y=-4x+12 /-3 /-3 /-3 y=-4/3x-4

HanAkoSolo (jamierox4ev3r):

4x-3(-4/3x-4)=12 4x-4x-4=12 +4 +4 0=8 x=0

HanAkoSolo (jamierox4ev3r):

this means that your final answer is a because it is -4 and 0

OpenStudy (anonymous):

no sorry i dont get it. @ jamiero4ev3r

HanAkoSolo (jamierox4ev3r):

first i solved for y, then i solved for x

OpenStudy (anonymous):

so i have to solve both?

OpenStudy (anonymous):

@Jamierox4ev3r

OpenStudy (zale101):

To find the y-intercept point, substitute 0 for the OTHER letter, x. and solve for the remaining letter y. Then make an ordered pair, putting that for the y-coordinate, and 0 for the x-coordinate.

OpenStudy (zale101):

Yes, you need to solve for both

OpenStudy (zale101):